# CryptoMoE: Privacy-Preserving and Scalable Mixture of Experts Inference via Balanced Expert Routing

### v1

@@ 5.4 Ablation Study
+Balanced Expert Routing 62.5 0.55
@@
Scalability. Our balanced expert routing strategy is scalable to larger models. Figure 9 shows the average accuracy of naive selection strategy (i.e., uniform random selection) and our CryptoMoE on Mixtral-13B/47B [48] and LLaMA4-Scout-17B/109B [1]. It can be seen that CryptoMoE consistently outperforms naive selection across all configurations. CryptoMoE maintains 100% accuracy on Mixtral even with $t=1.0$. On LLaMA4-Scout-109B, our CryptoMoEt=2.0 maintains 98.8% accuracy of the original model.
@@ Appendix A Expert activation patterns
Figure 10 shows the expert activation patterns in the 10th layer of DeepSeekMoE-16B [8] when processing text understanding versus mathematical reasoning tasks. Text understanding tasks include SIQA [13], OBQA [17], BoolQ [11], ARC-easy, ARC-challenge [16], HellaSwag [14], PIQA [12] and WinoGrande [15]. Mathematical reasoning tasks comprise GSM8K [9] and MathQA [10]. Expert #2 is disproportionately activated for text understanding, while Experts #3 and #60 show higher activation for mathematical reasoning.
@@ Appendix D Complexity Analysis of Batch MatMul Packing
In the original BOLT [24] packing method, each matrix $A_{i}$ is packed column-wise, with a single ciphertext holding $\frac{N}{t}$ columns. To accumulate results across columns, $\frac{N}{t}-1=O\left(\frac{N}{t}\right)$ rotations are required per ciphertext. For all $n$ matrix multiplications, the total number of ciphertexts is $O(\frac{ntd_{1}}{N})$, leading to an overall rotation cost of $O\left(\frac{N}{t}\cdot\frac{ntd_{1}}{N}\right)=O\left(nd_{1}\right)$.
@@
Integrate Baby-step Giant-step (BSGS) strategy into Batch MatMul packing. The BSGS algorithm is commonly used for ciphertext-plaintext MatMul to reduce HE rotations [24], decomposing rotations into local (baby-step) and global (giant-step) phases. This reduces the number of rotations per-MatMul from $O(d_{1})$ to $O\left(\sqrt{\frac{td_{1}d_{2}}{N}}\right)$. The total number of rotations is $O\left(n\sqrt{\frac{td_{1}d_{2}}{N}}\right)$. By combining BSGS with our batch packing, the effective parallel token dimension increases from $t$ to $nt$, since columns from all $n$ matrices are processed in parallel. This reduces the overall rotation complexity to $O\left(\sqrt{\frac{ntd_{1}d_{2}}{N}}\right)$, achieving a $\sqrt{n}\ \times$ improvement compared to applying BSGS independently to each matrix multiplication.

### v3

@@ 5.4 Ablation Study
+Balanced Expert Routing 50.9 0.55
@@
Scalability. Our balanced expert routing strategy is scalable to larger models. Figure 9 shows the average accuracy of naive selection strategy (i.e., uniform random selection) and our CryptoMoE on Mixtral-13B/47B [Jiang2024MixtralOE] and LLaMA4-Scout-17B/109B [LLAMA4]. It can be seen that CryptoMoE consistently outperforms naive selection across all configurations. CryptoMoE maintains 100% accuracy on Mixtral even with $t=1.0$. On LLaMA4-Scout-109B, our CryptoMoEt=2.0 maintains 98.8% accuracy of the original model.
@@ Appendix A Expert activation patterns
Figure 10 shows the expert activation patterns in the 10th layer of DeepSeekMoE-16B [dai2024deepseekmoe] when processing text understanding versus mathematical reasoning tasks. Text understanding tasks include SIQA [sap2019socialiqa], OBQA [OBQA], BoolQ [clark2019boolq], ARC-easy, ARC-challenge [arceasy_challenge], HellaSwag [zellers2019hellaswag], PIQA [bisk2020piqa] and WinoGrande [sakaguchi2021winogrande]. Mathematical reasoning tasks comprise GSM8K [gsm8k] and MathQA [amini2019mathqa]. Expert #2 is disproportionately activated for text understanding, while Experts #3 and #60 show higher activation for mathematical reasoning.
@@ Appendix D Complexity Analysis of Batch MatMul Packing
In the original BOLT [pang2023bolt] packing method, each matrix $A_{i}$ is packed column-wise, with a single ciphertext holding $\frac{N}{t}$ columns. To accumulate results across columns, $\frac{N}{t}-1=O\left(\frac{N}{t}\right)$ rotations are required per ciphertext. For all $n$ matrix multiplications, the total number of ciphertexts is $O(\frac{ntd_{1}}{N})$, leading to an overall rotation cost of $O\left(\frac{N}{t}\cdot\frac{ntd_{1}}{N}\right)=O\left(nd_{1}\right)$.
@@
Integrate Baby-step Giant-step (BSGS) strategy into Batch MatMul packing. The BSGS algorithm is commonly used for ciphertext-plaintext MatMul to reduce HE rotations [pang2023bolt], decomposing rotations into local (baby-step) and global (giant-step) phases. This reduces the number of rotations per-MatMul from $O(d_{1})$ to $O\left(\sqrt{\frac{td_{1}d_{2}}{N}}\right)$. The total number of rotations is $O\left(n\sqrt{\frac{td_{1}d_{2}}{N}}\right)$. By combining BSGS with our batch packing, the effective parallel token dimension increases from $t$ to $nt$, since columns from all $n$ matrices are processed in parallel. This reduces the overall rotation complexity to $O\left(\sqrt{\frac{ntd_{1}d_{2}}{N}}\right)$, achieving a $\sqrt{n}\ \times$ improvement compared to applying BSGS independently to each matrix multiplication.